You have given a function λ : R→R with the following properties (x ∈R, n ∈N): λ(n) = 0 , λ(x + 1) = λ(x) , λ(n + 1/ 2)= 1 Find two functions p,q : R→R with q(x) (is not equal to zero)= 0 for all x such that λ(x) = q(x)(p(x) + 1).
1
Expert's answer
2020-11-19T13:41:31-0500

We set "\lambda=sin(2\pi\,x)", "x\neq n+\frac12", "n\in{\mathbb{N}}" . For "x=n+\frac12" we put "\lambda=1". Then we put "q(x)=1," "p(x)=sin(2\pi x)-1," "x\neq n+\frac12,n\in{\mathbb{N}}" . For "x=n+\frac12" we set "p(x)=1".